OpenStudy (anonymous):

Convert 5 times pi divided by 6 radians into degrees I know what formula to put it in but don't know how to solve it could anyone help??

OpenStudy (phi):

you do this: \[ \frac{5 \pi}{6} \cdot \frac{180}{\pi} \] the \(\pi \)'s cancel. so type this into your calculator: \[ \frac{5 \cdot 180}{6} \] the answer will by the angle in degrees